OpenStudy (anonymous):

what is 5 and 1/4 divided by 1 and 5/9

OpenStudy (anonymous):

\[5 \frac{ 1 }{ 4 } \div 1\frac{ 5 }{ 9 }\]

OpenStudy (anonymous):

ya

OpenStudy (anonymous):

When you divide a fraction you change the division symbol to a multiplication and flip the second fraction\[5 \frac{ 1 }{ 4 } \times 1\frac{ 9 }{ 4 }\]

OpenStudy (anonymous):

Make more sense?